# Junhan Kim    Ho-young Kim    Eulrang Cho    Chungman Lee    Joonyoung Kim    Yongkweon Jeon

### v2

@@ 3.3 Efficient Implementation of BoA
    Results for high bit-widths and results on LLaMA1 and OPT models are provided in Section C.1 due to the page limitation.
@@
    Results for high bit-widths and results on LLaMA1 models are provided in Section C.2 due to the page limitation.
4 Experiments
4.1 Experimental Setup
We conduct experiments on OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023a)), LLaMA2 (Touvron et al., 2023b), and LLaMA3. As a calibration dataset, we use 128 random sequences of length 2048 from WikiText-2 (Merity et al., 2016), as in (Shao et al., 2023; Ma et al., 2024; Lin et al., 2024a; Ashkboos et al., 2024; Liu et al., 2024). We report the perplexity (PPL) score on the WikiText-2 test dataset and accuracy on eight zero-shot commonsense reasoning tasks (ARC-challenge (Arc-c) and ARC-easy (Arc-e) (Clark et al., 2018), BoolQ (BQ) (Clark et al., 2019), HellaSwag (HS) (Zellers et al., 2019), LAMBADA (LAMB) (Paperno et al., 2016), OpenbookQA (OBQA) (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), and WinoGrande (WG) (Sakaguchi et al., 2021)). All experiments were conducted using a single NVIDIA H100 GPU (80 GB).
4.2 Weight-Only Quantization Results
Performance of integer weight optimization
We first evaluate the weight-only quantization performance of the proposed BoA. In this experiment, to solely compare the performance of the integer weight optimization, we do not apply any transform to the models, such as smoothing (Xiao et al., 2023; Lin et al., 2024b; Jeon et al., 2023; Shao et al., 2023; Ma et al., 2024), rotation (Ashkboos et al., 2024; Liu et al., 2024), or permutation (Lin et al., 2024a).
In Table 3 and Tables 7-9 (see Section C.1), we summarize the PPL and zero-shot task performances of BoA. For comparison, we include the performance of the conventional GPTQ and the rounding-to-nearest (RTN) method which naively assigns the nearest quantization bin. We observe BoA and GPTQ exhibit reasonable PPL even for INT2 quantization where RTN collapses significantly (i.e., PPL is almost $10^{4}$). This is because BoA and GPTQ minimize the task loss degradation rather than the weight quantization error $\Delta\mathbf{W}$. As evident, BoA significantly surpasses GPTQ on all models in both PPL and zero-shot accuracy. For example, BoA achieves 10%p accuracy improvement on the 3-bit quantized LLaMA2-13B and LLaMA3.2-1B models. In addition, BoA shows 20%p higher accuracy for the 3-bit quantized LLaMA3.2-3B model. The key factor leading to such an outstanding performance is that BoA considers inter-layer dependencies within the attention module by exploiting the proposed attention-aware Hessians while GPTQ assumes the independence of layers.
@@ 4.2 Weight-Only Quantization Results
To improve the quantization performance, recent studies have transformed models by applying smoothing, rotation, or permutation (Shao et al., 2023; Ma et al., 2024; Ashkboos et al., 2024; Liu et al., 2024; Lin et al., 2024a). In this experiment, we evaluate BoA under those model transformations (see Table 4 and Tables 10-11 in Section C.2). For conventional methods, we ran the official codes provided by the authors and reported the obtained results; details for implementing conventional methods are provided in Section C.2. When measuring the performance of BoA, we also transformed the model for a fair comparison. We applied QuaRot for the transformation because QuaRot does not require any training and incurs no extra costs during the actual inference (Ashkboos et al., 2024).
We observe that the performance of the proposed BoA is boosted significantly by applying the transformation. For example, the PPLs of the 2-bit quantized LLaMA2-13B / LLaMA3-8B models improve from 18.33 / 21.70 to 8.237 / 15.24, respectively (see Tables 3 and 4). For all models, we achieve at least 5%p improvement in the zero-shot accuracy (in particular 11%p improvement for the 2-bit quantized LLaMA2-13B model). We note that the performance gap between the proposed BoA and GPTQ remains significant after the model transformation; for example, the zero-shot accuracy of BoA on the 2-bit quantized LLaMA2-7B model is 12%p larger than that obtained by QuaRot-GPTQ.
Table 5: Processing time (hour) of different quantization methods considering inter-layer dependencies
Method Optimization Type LLaMA Model Size
7B 13B 30B
OmniQuant Gradient-based 1.827 3.322 7.661
AffineQuant 4.323 8.406 21.37
DuQuant + LWC 1.221 2.077 4.547
BoA One-shot 0.961 1.553 3.295
@@
    Results for other configurations are provided in Section C.3 due to the page limitation.
@@ 4.3 Weight-Activation Quantization Results
We now evaluate the weight-activation quantization performance (see Table 6 and Tables 12-14 in Section C.3). As in previous studies (Shao et al., 2023; Ma et al., 2024; Lin et al., 2024a; Ashkboos et al., 2024; Liu et al., 2024), we apply per-token nearest-rounding quantization for input activations and KV caches. We use the notation ‘WxAyKVz’ to denote the x-bit weight quantization, y-bit input activation quantization, and z-bit KV cache quantization. When measuring the performance of RTN, GPTQ, and the proposed BoA, we use SpinQuant for the model transformation as SpinQuant outperforms QuaRot by optimizing the transformation for the activation quantization (Liu et al., 2024).
From Table 6 and Tables 12-14 in Section C.3, we observe that the outstanding weight-quantization performance of BoA leads to state-of-the-art performance for the weight-activation quantization. In particular, the performance gain is noticeable for low-bit (e.g., W2A4KV4 and W2A4KV16). For example, BoA achieves 10%p accuracy improvement on LLaMA2-7B and 12.5%p improvement on LLaMA2-13B over SpinQuant-GPTQ (see Table 6). Furthermore, compared to OmniQuant, AffineQuant, and DuQuant, BoA achieves 8%p, 13%p, 13%p, 5%p, and 7%p improvement on LLaMA2-7B, LLaMA2-13B, LLaMA3-8B, LLaMA3.2-1B, and LLaMA3.2-3B, respectively.
@@ 5 Conclusion
In this paper, we proposed a novel backpropagation-free PTQ algorithm called BoA. To consider the inter-layer dependencies within the attention module, we approximated the Hessian matrices by exploiting the attention reconstruction error rather than the layer-wise reconstruction error. Furthermore, to mitigate the computational overhead incurred by the proposed attention-aware Hessians, we incorporated several techniques including Hessian relaxation, efficient computation of inverse Hessians, and head-wise simultaneous quantization. Finally, from extensive experiments, we demonstrated the efficacy of the proposed BoA.
@@ Appendix A Proof of Footnote 2
where $\Delta\mathbf{w}=\operatornamewithlimits{vec}(\Delta\mathbf{W})$. In addition, by 13b and 13c, we have
@@
where (a) follows from 13b. This completes the proof.
Appendix B Refined Weight-update Formula
@@
By taking similar steps as above, we can easily generalize 14 for the $j$-th row as follows:
@@
Appendix C Additional Experimental Results
C.1 Performance of Weight-Rounding Optimization
@@
C.2 Comparison with Transformation-based Methods
We compare the weight-only quantization performances of the proposed BoA and existing transformation-based methods. For conventional methods, we ran the official codes provided by the authors and reported the obtained results. For OmniQuant (Shao et al., 2023) and AffineQuant (Ma et al., 2024), we activated both learnable equivalent transformation (LET) and learnable weight clipping (LWC) options. For AffineQuant, we did not activate the use-ln-matrix option because this adds extra affine transformation between the normalization layer and linear layers, which incurs additional processing time during the inference. QuaRot-RTN and QuaRot-GPTQ mean that RTN and GPTQ have been applied for the weight quantization after transforming models via QuaRot (Ashkboos et al., 2024). For DuQuant, we report the results obtained with and without activating the LWC option as in the original paper (Lin et al., 2024a). When measuring the performance of the proposed BoA, we also transformed the model for fair comparison. We applied QuaRot for the transformation because QuaRot does not require training and incurs no extra costs during the actual inference (Ashkboos et al., 2024).
@@ C.2 Comparison with Transformation-based Methods
C.3 Weight-Activation Quantization Results
We evaluate the weight-activation quantization performances of the proposed BoA. As in prior works (Shao et al., 2023; Ma et al., 2024; Lin et al., 2024a; Ashkboos et al., 2024; Liu et al., 2024), we apply per-token nearest-rounding quantization for input activations and KV caches. We use the notation ‘WxAyKVz’ to denote the x-bit weight quantization, y-bit input activation quantization, and z-bit KV cache quantization. When measuring the performance of RTN, GPTQ, and the proposed BoA, we use SpinQuant for the model transformation as SpinQuant outperforms QuaRot by optimizing the transformation for the activation quantization (Liu et al., 2024). When optimizing the rotation matrix in SpinQuant, we quantize both weights and activations for SpinQuant-RTN and quantize only activations for SpinQuant-GPTQ and the proposed BoA, as in the original paper (Liu et al., 2024). In this appendix, we provide the results on LLaMA3 models for various quantization configurations.

### v3

@@ 3.3 Efficient Implementation of BoA
    Results for high bit-widths and results on LLaMA1 and OPT models are provided in Section D.1 due to the page limitation.
We measure quantization processing times of BoA with and without simultaneous quantization to evaluate how much the head-wise joint quantization can accelerate the quantization process. From Table 2, we observe that BoA requires a significantly long processing time without the simultaneous quantization (more than one day even for the 7B model). This is because all rows need to be quantized sequentially (e.g., 4096 rows are quantized sequentially for LLaMA-7B) and thus the massive compute capabilities of modern GPUs cannot be utilized properly. As evident, we can significantly reduce the processing time by applying the head-wise joint quantization (more than 40 times reduction on the 30B model).
4 Experiments
4.1 Experimental Setup
We conduct experiments on OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023a)), LLaMA2 (Touvron et al., 2023b), and LLaMA3. As in previous studies (Shao et al., 2023; Ma et al., 2024; Lin et al., 2024; Ashkboos et al., 2024; Liu et al., 2024), we construct a calibration dataset by sampling 128 random sequences of length 2048 from WikiText-2 (Merity et al., 2016). As a performance metric, we use the perplexity (PPL) score on the WikiText-2 test dataset and accuracy on eight zero-shot commonsense reasoning tasks (ARC-challenge (Arc-c) and ARC-easy (Arc-e) (Clark et al., 2018), BoolQ (BQ) (Clark et al., 2019), HellaSwag (HS) (Zellers et al., 2019), LAMBADA (LAMB) (Paperno et al., 2016), OpenbookQA (OBQA) (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), and WinoGrande (WG) (Sakaguchi et al., 2021)). All experiments were conducted using a single NVIDIA H100 GPU (80 GB).
When determining a quantization order in BoA, the heuristic introduced by GPTQ can be used; the column/row corresponding to the largest $\text{diag}(\mathbf{H}_{\text{col}})/\text{diag}(\mathbf{H}_{\text{row}})$ (i.e., the most quantization-sensitive column/row) is first quantized for better compensation. Empirically, we observed that this heuristic could occasionally enhance the performance, yet at other times, it may result in inferior performance. For GPTQ and the proposed BoA, we conduct experiments with and without this heuristic and report the better results.
4.2 Weight-Only Quantization Results
Comparison with GPTQ
We first compare the proposed BoA with GPTQ to demonstrate the efficacy of the proposed attention-aware Hessians. In this experiment, we do not apply any model transformation method (e.g. QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2024)) to solely evaluate the effectiveness of the proposed Hessians.
In Table 3 and Tables 8-10 (see Section D.1), we summarize the PPL and zero-shot task performances of BoA and GPTQ. We also include the performance of the rounding-to-nearest (RTN) method which naively assigns the nearest quantization bin. We observe BoA and GPTQ exhibit reasonable PPL even for INT2 quantization where RTN collapses significantly (i.e., PPL is almost $10^{4}$). This is because BoA and GPTQ minimize the task loss degradation rather than the weight quantization error $\Delta\mathbf{W}$. As evident, BoA significantly surpasses GPTQ on all models in both PPL and zero-shot accuracy. For example, BoA achieves 10%p accuracy improvement on the 3-bit quantized LLaMA2-13B and LLaMA3.2-1B models. In addition, BoA shows 20%p higher accuracy for the 3-bit quantized LLaMA3.2-3B model. The key factor leading to such an outstanding performance is that BoA considers inter-layer dependencies within the attention module by exploiting the proposed attention-aware Hessians while GPTQ assumes the independence of layers.
Table 4: Memory costs of GPTQ and the proposed BoA
Method Memory Cost (GB) Wiki2 PPL ($\downarrow$)
7B 13B 30B 7B 13B 30B
GPTQ 4.426 5.872 8.271 19.63 34.63 9.770
BoA 9.550 16.56 32.79 10.28 8.309 6.669
Relaxed BoA 6.446 8.397 11.55 10.53 8.700 7.007
    Attention-aware Hessians have been applied to query and key, but not to value.
In Table 4, we summarize the memory costs of BoA and GPTQ. We observe that BoA requires larger memory because BoA additionally uses outputs of other layers to consider inter-layer dependencies. It is worth mentioning that when the memory resource is limited, BoA can still be used with a slight relaxation. Noting that the large memory cost of BoA is attributable to the Hessian for the value projection ($\mathbf{H}_{\text{col},h}=2\mathbf{X}\mathbf{A}_{h}^{T}\mathbf{A}_{h}\mathbf{X}^{T}$ in 9) whose shape is $H\times d\times d$, we can greatly reduce the memory cost by exploiting the standard Hessian in 3 for the value projection and applying the proposed attention-aware Hessians only for query and key projections. In doing so, BoA needs slightly more memory (e.g. 3 GB for 30B; see Table 4) yet still performs much better than GPTQ.
We now compare the processing times of BoA and GPTQ. As evident from Table 5, BoA requires a longer processing time than that needed by GPTQ. This is because GPTQ quantizes all the rows simultaneously, while BoA sequentially quantizes them to compensate for the quantization error of each row (see Figure 1(b)). Clearly, there exists a trade-off between quantization speed and accuracy. In real cases, when one needs to consider inter-layer dependencies to preserve the performance of the original model as much as possible, the proposed BoA would be an intriguing solution when compared to existing gradient-based approaches (see Table 5 and Table 6). When faster quantization is required, one possible solution is to reduce the number of sequential quantizations by quantizing multiple rows in each head simultaneously, which will be considered in our future studies.
Table 5: Processing time (hour) of different quantization methods
Method Inter-layer Dependency Optimization Type LLaMA Model Size
7B 13B 30B
GPTQ X one-shot 0.12 0.20 0.43
OmniQuant O gradient -based 1.83 3.32 7.66
AffineQuant 4.32 8.41 21.4
DuQuant+LWC 1.22 2.08 4.55
BoA O one-shot 0.96 1.55 3.30
@@ 4.2 Weight-Only Quantization Results
    Results for high bit-widths and results on LLaMA1 models are provided in Section D.2 due to the page limitation.
@@
To improve the quantization performance, recent studies have transformed models by applying smoothing, rotation, or permutation (Shao et al., 2023; Ma et al., 2024; Ashkboos et al., 2024; Liu et al., 2024; Lin et al., 2024). In this experiment, we evaluate BoA under those model transformations (see Table 6 and Tables 11-12 in Section D.2). For conventional methods, we ran the official codes provided by the authors and reported the obtained results; details for implementing conventional methods are provided in Section D.2. When measuring the performance of BoA, we also transformed the model for a fair comparison. We applied QuaRot for the transformation because QuaRot does not require any training and incurs no extra costs during the actual inference (Ashkboos et al., 2024).
We observe that the performance of the proposed BoA is boosted significantly when applying the transformation. For example, the PPLs of the 2-bit quantized LLaMA2-13B / LLaMA3-8B models improve from 18.33 / 21.70 to 8.237 / 15.24, respectively (see Tables 3 and 6). For INT3 quantization, BoA almost preserves the performance of the original full-precision model; the accuracy drop is 2.3%p for LLaMA3-8B and 1.3%p for LLaMA2-13B (see Table 11). For all models, we achieve at least 5%p improvement in the zero-shot accuracy (in particular 11%p improvement for the 2-bit quantized LLaMA2-13B model). We note that the performance gap between BoA and GPTQ remains significant after the model transformation; for example, the zero-shot accuracy of BoA on the 2-bit quantized LLaMA2-7B model is 12%p larger than that obtained by QuaRot-GPTQ.
@@
    Results for other configurations are provided in Section D.3 due to the page limitation.
@@ 4.3 Weight-Activation Quantization Results
We now evaluate the weight-activation quantization performance (see Table 7 and Tables 13-15 in Section D.3). As in previous studies (Shao et al., 2023; Ma et al., 2024; Lin et al., 2024; Ashkboos et al., 2024; Liu et al., 2024), we quantize input activations to all linear layers and KV caches with the Min-Max nearest-rounding quantizer where quantization parameters are determined dynamically for each token. We use the notation ‘WxAyKVz’ to denote the x-bit weight quantization, y-bit input activation quantization, and z-bit KV cache quantization. In this experiment, when measuring the performance of RTN, GPTQ, and the proposed BoA, we use SpinQuant (instead of QuaRot) for the model transformation. It is worth noting that while QuaRot uses Hadamard matrices (that are independent of data) for the rotation (Ashkboos et al., 2024), SpinQuant optimizes rotation matrices that make models more robust to the activation quantization and thus could outperform QuaRot (Liu et al., 2024). In our experiments, rotation matrices have been optimized with respect to activation-only quantized networks (i.e., weights are fixed with full-precision) as in (Liu et al., 2024).
From Table 7 and Tables 13-15 in Section D.3, we observe that the outstanding weight-quantization performance of the proposed BoA leads to the state-of-the-art performance for the weight-activation quantization. In particular, the performance gain is noticeable for low-bit (e.g., W2A4KV4 and W2A4KV16). For example, compared to SpinQuant-GPTQ, BoA achieves 10%p accuracy improvement on LLaMA2-7B and 12.5%p improvement on LLaMA2-13B (see Table 7). Compared to the conventional gradient-based methods (i.e. OmniQuant, AffineQuant, and DuQuant), BoA achieves 8%p, 13%p, 13%p, 5%p, and 7%p improvement on LLaMA2-7B, LLaMA2-13B, LLaMA3-8B, LLaMA3.2-1B, and LLaMA3.2-3B, respectively.
@@ 5 Conclusion
In this paper, we proposed a novel backpropagation-free PTQ algorithm called BoA. To consider the inter-layer dependencies within the attention module, we approximated the Hessian matrices by exploiting the attention reconstruction error rather than the layer-wise reconstruction error. To mitigate the computational overhead incurred by the proposed attention-aware Hessians, we also incorporated several techniques, including Hessian relaxation, efficient computation of inverse and Cholesky decomposition of attention-aware Hessians, and simultaneous quantization of different attention heads. Finally, from extensive experiments, we demonstrated the efficacy of the proposed BoA.
Acknowledgment
We would like thanks to Daehyun Kim, Ph.D., Hyeonmok Ko, Ph.D., and Hyemi Jang, Ph.D. for their helpful discussion.
@@ Appendix A Proof of Footnote 2
where $\Delta\mathbf{w}=\operatornamewithlimits{vec}(\Delta\mathbf{W})$. In addition, by 17b and 17c, we have
@@
where (a) follows from 17b. This completes the proof.
Appendix B Attention-aware Hessians for Models Exploiting RoPE
As mentioned, when RoPE is applied, the proposed surrogate $\|\mathbf{K}_{h}\Delta\mathbf{Q}_{h}^{T}\|_{F}^{2}$ used to develop the attention-aware Hessian $\mathbf{H}^{(\mathbf{w}_{Q,h})}$ in 11 changes to $\|\widetilde{\mathbf{K}}_{h}\Delta\widetilde{\mathbf{Q}}_{h}^{T}\|_{F}^{2}$, where $\widetilde{\mathbf{K}}_{h}=\operatornamewithlimits{RoPE}(\mathbf{K}_{h})$ and $\widetilde{\mathbf{Q}}_{h}=\operatornamewithlimits{RoPE}(\mathbf{Q}_{h})$. Let $\mathbf{R}_{\ell}$ be the rotary matrix for the $\ell$-th token (see eq. (15) in (Su et al., 2023)) and $\widetilde{\mathbf{Q}}_{h}^{T}=[\widetilde{\mathbf{q}}_{h,1}\ \ldots\widetilde{\mathbf{q}}_{h,L}]$, then the objective can be expressed as
$\|\widetilde{\mathbf{K}}_{h}\Delta\widetilde{\mathbf{Q}}_{h}^{T}\|_{F}^{2}=\sum_{\ell=1}^{L}\|\widetilde{\mathbf{K}}_{h}\Delta\widetilde{\mathbf{q}}_{h,\ell}\|_{2}^{2}=\sum_{\ell=1}^{L}\|\widetilde{\mathbf{K}}_{h}\Delta(\mathbf{R}_{\ell}\mathbf{W}_{Q,h}\mathbf{x}_{\ell})\|_{2}^{2}=\sum_{\ell=1}^{L}\|\widetilde{\mathbf{K}}_{h}\mathbf{R}_{\ell}\Delta\mathbf{W}_{Q,h}\mathbf{x}_{\ell}\|_{2}^{2},$
which yields the following Hessian equation (see Footnote 2):
$\mathbf{H}^{(\mathbf{w}_{Q,h})}=\sum_{\ell=1}^{L}(2\mathbf{x}_{\ell}\mathbf{x}_{\ell}^{T}\otimes\mathbf{R}_{\ell}^{T}\widetilde{\mathbf{K}}_{h}^{T}\widetilde{\mathbf{K}}_{h}\mathbf{R}_{\ell}).$
Finally, we take the factorized approximation for the summation of Kronecker products (i.e., $\mathbb{E}[\mathbf{M}_{1}\otimes\mathbf{M}_{2}]\approx\mathbb{E}[\mathbf{M}_{1}]\otimes\mathbb{E}[\mathbf{M}_{2}]$; see eq. (20) in (Botev et al., 2017)):
$\mathbf{H}^{(\mathbf{w}_{Q,h})}\approx\sum_{\ell=1}^{L}2\mathbf{x}_{\ell}\mathbf{x}_{\ell}^{T}\otimes\frac{1}{L}\sum_{\ell=1}^{L}\mathbf{R}_{\ell}^{T}\widetilde{\mathbf{K}}_{h}^{T}\widetilde{\mathbf{K}}_{h}\mathbf{R}_{\ell}=2\mathbf{X}\mathbf{X}^{T}\otimes\frac{1}{L}\sum_{\ell=1}^{L}\mathbf{R}_{\ell}^{T}\widetilde{\mathbf{K}}_{h}^{T}\widetilde{\mathbf{K}}_{h}\mathbf{R}_{\ell}.$
By taking similar steps, the attention-aware Hessian for the key projection weight $\mathbf{W}_{K,h}$ with RoPE can be established as
$\mathbf{H}^{(\mathbf{w}_{K,h})}=2\mathbf{X}\mathbf{X}^{T}\otimes\frac{1}{L}\sum_{\ell=1}^{L}\mathbf{R}_{\ell}^{T}\widetilde{\mathbf{Q}}_{h}^{T}\widetilde{\mathbf{Q}}_{h}\mathbf{R}_{\ell}.$
Appendix C Refined Weight-update Formula
@@
By taking similar steps as above, we can easily generalize 18 for the $j$-th row as follows:
@@
Appendix D Additional Experimental Results
D.1 Performance of Weight-Rounding Optimization
@@
D.2 Comparison with Transformation-based Methods
We compare the weight-only quantization performances of the proposed BoA and existing transformation-based methods. For conventional methods, we ran the official codes provided by the authors and reported the obtained results. For OmniQuant (Shao et al., 2023) and AffineQuant (Ma et al., 2024), we activated both learnable equivalent transformation (LET) and learnable weight clipping (LWC) options. For AffineQuant, we did not activate the use-ln-matrix option because this adds extra affine transformation between the normalization layer and linear layers, which incurs additional processing time during the inference. QuaRot-RTN and QuaRot-GPTQ mean that RTN and GPTQ have been applied for the weight quantization after transforming models via QuaRot (Ashkboos et al., 2024). For DuQuant, we report the results obtained with and without activating the LWC option as in the original paper (Lin et al., 2024). When measuring the performance of the proposed BoA, we also transformed the model for fair comparison. We applied QuaRot for the transformation because QuaRot does not require training and incurs no extra costs during the actual inference (Ashkboos et al., 2024).
@@ D.2 Comparison with Transformation-based Methods
D.3 Weight-Activation Quantization Results
We evaluate the weight-activation quantization performances of the proposed BoA. As in prior works (Shao et al., 2023; Ma et al., 2024; Lin et al., 2024; Ashkboos et al., 2024; Liu et al., 2024), we apply per-token nearest-rounding quantization for input activations and KV caches. We use the notation ‘WxAyKVz’ to denote the x-bit weight quantization, y-bit input activation quantization, and z-bit KV cache quantization. When measuring the performance of RTN, GPTQ, and the proposed BoA, we use SpinQuant for the model transformation as SpinQuant outperforms QuaRot by optimizing the transformation for the activation quantization (Liu et al., 2024). When optimizing the rotation matrix in SpinQuant, we quantize both weights and activations for SpinQuant-RTN and quantize only activations for SpinQuant-GPTQ and the proposed BoA, as in the original paper (Liu et al., 2024). In this appendix, we provide the results on LLaMA3 models for various quantization configurations.